# An Adaptive Online Smoother with Closed-Form Solutions and Information-Theoretic Lag Selection for Conditional Gaussian Nonlinear Systems

### v1

@@ 4 Applications of the Adaptive-Lag Online Smoother
This section demonstrates the application of the online smoother in addressing three key scientific challenges: (a) state estimation and causality detection, especially in the presence of intermittency and extreme events, (b) evaluating computational and storage efficiency in a high-dimensional LDA application, and (c) developing an online parameter estimation algorithm with partial observations.
The simulations for all the numerical experiments in this work were carried out on an AMD RyzenTM 7 5800H mobile CPU of 8 cores, 16 threads, with an average clock speed of 3.2 GHz and turbo boost up to 4.4 GHz, with a TDP of 45W. In terms of memory, a two-memory-module configuration was used, with 2x16GB DDR4 RAM sticks at 3200MT/s.
@@ 4.1 Detecting Causal Evidence Using the Adaptive-Lag Online Smoother
${\,\rm d}u(t)$ $=(-d_{u}u(t)+\gamma u(t)v(t)+F_{u}){\,\rm d}t+\sigma_{u}{\,\rm d}W_{u}(t),$ (4.1)
${\,\rm d}v(t)$ $=(-d_{v}v(t)-\gamma u(t)^{2}+F_{v}){\,\rm d}t+\sigma_{vu}u(t){\,\rm d}W_{u}(t)+\sigma_{v}{\,\rm d}W_{v}(t),$ (4.2)
where all parameters are constants, with $d_{u},d_{v}\in\mathbb{R}$ and $\gamma,\sigma_{u},\sigma_{vu},\sigma_{v}>0$, while $W_{u}$ and $W_{v}$ are two mutually independent Wiener processes. In (4.1)–(4.2), The variable $u$ represents an observed or resolved mode in the turbulent signal, interacting with the unresolved mode $v$ through quadratic nonlinearities. As $v$ influences the damping of $u$, variations in $v$ cause the time series of $u$ to display intermittency and extreme events. The system (4.1)–(4.2) fits exactly the CGNS framework in (2.1)–(2.2), with $\mathbf{x}=x:=u$ and $\mathbf{y}=y:=v$. The following parameter values are used for this numerical case study:
@@
The parameters are assigned to induce intermittent and extreme events in the observable variable $u$ through the stochastic damping generated by the unobserved variable $v$, with the goal of producing highly non-Gaussian PDFs for both observable and unobserved dynamics. These values are also chosen to address potential issues of observability [13, 115, 116, 117, 118], which is linked to causal detection using the adaptive online smoother. The coupled system in (4.1)–(4.2) loses practical observability when the observed process $u$ provides no information about the unobserved variable $v$. Intuitively, this occurs during phases where $u\approx 0$, rendering $v$ ineffective in contributing to the dynamics of $u$. A numerical integration time step of $\Delta t=0.005$ is used, which also defines the uniform rate of obtaining the observations for the online smoother, with a total simulation time of $T=60$ units.
@@
Figure 4.1 presents various plots comparing the filter and smoother posterior estimated states at different levels. Panel (a) displays the trajectory of the observed variable $u$, while panel (b) displays that for the unobserved variable $v$. For the latter, we also include the posterior means calculated using Theorems 2.2 and 2.3, along with the first two standard deviations away from the mean, represented by accordingly colored shaded regions. The posterior estimated states overall follow the true values with relatively low uncertainty. As expected, the smoother provides a more accurate estimation during the uncertain quiescent periods of the unobserved variable. During the extreme events which are characterized by a high signal-to-noise ratio, both methods lead to smaller uncertainty. Panel (c) plots the information gain of the filter and smoother posterior distributions relative to the prior equilibrium distribution of $v$ on a logarithmically scaled y-axis. Since both the equilibrium (computed numerically using many long simulated trajectories; not shown here) and posterior distributions are Gaussian, the information gain in panel (c) is calculated using the signal-dispersion decomposition formula of relative entropy in (3.26). The temporal evolution of relative entropy reveals that the information gain from the smoother posterior Gaussian distribution is almost everywhere larger than that from the filter solution, as is expected. Notably, the information gain corresponding to the onset of the intermittent phases of $u$ shoots up due to the strong signal-to-noise ratios. Note that the information gain using the smoother is more significant than the filter at these phases and the more uncertain quiescent phases due to additional observational information. Panel (d) shows the PDF of $u$ generated by its complete signal over $t\in[0,60]$, alongside the corresponding Gaussian fit, i.e., the normal distribution with the same mean and standard deviation as the truth over the associated period. Due to the intermittent extreme events, we can discern that the induced density is highly non-Gaussian, with an apparent positive skewness and a heavy tail. Panel (e) is the same as (d) for $v$, which also includes the densities corresponding to the filter and smoother solutions. We can again see that the density corresponding to $v$ displays some non-Gaussian features. Furthermore, as is expected, the PDF of the smoother solution is better equipped to approximate that of the truth compared to the filter one by better resolving the tail behavior and first few ordered moments.
@@
Figure 4.1: Panel (a): Trajectory of the observable variable $u$. Panel (b): True trajectory of the unobserved variable $v$ (in blue), alongside the filter (in green) and smoother (in red) posterior mean time series. Their respective first two standard deviations away from the mean state are also plotted through correspondingly colored shaded regions, with the standard deviation in this case being exactly equal to the square root of the respective posterior covariance. The threshold above which $v$ acts as anti-damping to $u$, i.e., the line $y=d_{u}/\gamma=1/6$, is also plotted in a dashed black line. Panel (c): Temporal evolution of the information gain of the filter and smoother posterior Gaussian statistics beyond the statistical attractor of the unobserved variable. This is calculated using the signal-dispersion decomposition of the relative entropy in (3.26). A logarithmic scale is used for the y-axis. Panel (d): PDF of $u$ calculated using the observations over $t\in[0,60]$ (in blue), alongside its Gaussian fit density defined by the mean and standard deviation of the signal in the same time period (in dashed black). Panel (e): Same as (d) but for the unobserved variable $v$, together with the PDFs corresponding to the filter (in green) and smoother (in red) posterior mean time series.
@@
Figure 4.2 focuses on a shorter intermittent period ($t\in[32.5,40]$) to assess the performance of the adaptive-lag online smoother, using the LSDF sequence, $\{\sigma^{j,n}\}_{R_{n}\leq j\leq n-1}$, with the hyperparameters set to $b=400$ (which corresponds to $b\Delta t=2$ simulation time units), $\delta=10^{-6}$, and using MATLAB’s default value for $w$ in stdfilt which is $w=3$. For this case study, equivalent results can be obtained when instead the entropy sequence itself, $\left\{\mathcal{P}\left(p^{j,n}_{\text{updated}},p^{j,n}_{\text{lagged}}\right)\right\}_{R_{n}\leq j\leq n-1}$, is being used, as already discussed in Section 3.4.2, by appropriately adjusting $\delta$. Panel (a) shows the signals of $u$ and $v$ over this period, while panels (b)–(f) illustrate how the online smoother adjusts its state estimation as new observations arrive at five different time instants during this period. The online smoother is compared against the filter solution, which can also produce real-time estimated states with the sequential arrival of observations. We focus on panels (c) and (d). When using the filter solution, the estimation of $v$ before $u$ reaches the peak of the extreme event contains significant errors. This is because the growth of $v$, once it exceeds the threshold of $d_{u}/\gamma=1/6$ to become anti-damping, triggers the extreme event in $u$ (the cause). However, the extreme event (the effect) does not appear instantly, thus introducing a delay in the onset of these instabilities. This is why at $t=36.50$ (panel (c)), without seeing the future information from the development of the extreme event, the causal variable $v$ cannot be accurately recovered by the filter. By extension, due to the data from the extreme event in the observed signal simply not being available yet at the current time instant, this, in a sense, contaminates the online smoother state, which faithfully follows the filter one regardless of how long the adaptive lag is. Nevertheless, at $t=37.00$ (panel (d)), with the observed trajectory currently undergoing the effects of the instability caused by $v$, this allows the adaptive online smoother, which utilizes future information for state estimation unlike the filter one, to correct its previous estimated states based on the value of the calculated adaptive lag, at most for $b\Delta t=2$ time units into the past. This is explicitly seen by the online smoother posterior mean, which now better approximates the true signal compared to its state half a time unit ago, thus demonstrating its property to systematically correct biases when future data becomes available. This attribute is crucial to help identify the current state, therefore, outperforming the filter solution. This crucial discrepancy between the filter and online smoother estimates reveals that information in the dynamics flows from the current state of the unobserved variable (cause) to the future state of the observed one (effect), highlighting a causal relationship between these states. Panel (g) presents the adaptive lag values generated during this period by using the LSDFs of the information gain sequence outlined in (3.31). The lag in the online smoother likewise reflects the aforementioned time delay in the causal relationship, with a peak in adaptive lag values at the onset of the extreme event, i.e., the causal period. These are essential to correct past errors and effectively recover the state of the cause $v$, which triggered the extreme event. In contrast, smaller lag values are needed during the effect period, specifically during the demise of this intermittent event, which is characterized by a high signal-to-noise ratio where both the filter and smoother estimates can effectively capture the relative deterministic behavior. As such, lower adaptive lag values are required since the “lagged” distribution in (3.29) can yield a quickly decaying information gain function as $j$ decreases from $n-1$ to $R_{n}$ (or to an already small sequence of information gains for these values of $n$, i.e., below the threshold parameter $\delta$ uniformly in $j$), thus leading to small values for $L_{n}$ in (3.31). After all, the lag essentially functions as an interpolation between the past state estimates and a “full” (since $b$ limits how much into the past we can correct) backward-run smoother solution of the online smoother. As such, a close-to-zero lag value signifies that more trust is put in the statistics from the previously calculated online posterior distributions since the observations obtained during this period have tiny impact regions.
In panel (h) of Figure 4.2, the LSDF criterion sequences used to define the adaptive lag in panel (g) (see (3.31)) at the time instants showed in panels (b) through (f) are plotted, on a logarithmically scaled y-axis. Although each one is defined for $R_{n}\leq j\leq n-1$, here these are plotted on a common x-axis for ease of comparison, with the lag on the x-axis also being measured in simulation time units for clarity, i.e., the x-axis is over $[0,b\Delta t]$. The tolerance value of $\delta=10^{-6}$ is also plotted in a black dashed line. Based on (3.31), $b\Delta t=2$ minus the first time on which each curve goes below this threshold line, defines the respective adaptive lag value (i.e., for the corresponding $n$) in panel (g). We note that the minute oscillations observed are by virtue of the LSDF sequence essentially functioning as the numerical first-order derivative of the information gains. Increasing the window size $w$ or using the sequence of relative entropies instead yields equivalent curves but without these numerical artifacts. Of significance are the curves that cover the cause and effect of the extreme event around $[36.5,37.5]$, specifically the curves for $t=37.00$ (in green) and $t=38.50$ (in magenta). The former only captures its cause, so the damping in the information gain criterion is only barely observed for $j$ close to $n-1$ (i.e., close to $b\Delta t=2$ in this common and absolute x-axis). On the other hand, for the curve corresponding to $t=38.50$, with $j$ decreasing from $n-1$ down to $R_{n}$ (i.e., from $b\Delta t=2$ down to 00), we can see that there is a significant change in the exponential behavior of the information gain criterion at around $t=37.50$ which is the demise point of the extreme event in $u$, with the high signal-to-noise ratio extreme event between $[36.5,37.5]$ inducing a regime switch in the information gain criterion to a smaller exponential function with roughly the same rate, through sub-unit damping, which in turn yields the significant drop in the gained information as a function of $j$. This has the ulterior outcome of smaller adaptive lag values during this period, as shown in panel (g), compared to the other periods. For the sake of completion, observe how this behavior also starts to form for the green curve corresponding to $t=37.00$ for $j$ close to $R_{n}$ (i.e., close to 00) since it captures the demise of the effect in the previous (and milder) extreme event observed at around $t=35$. In contrast, the blue curve for $t=40.00$ is almost a straight line in this semilog-y plot, without any regime switches, since it covers a quiescent period in the signal.
@@
Finally, in panel (i) of Figure 4.2, the spectral radii of the update matrices $\mathbf{D}^{j,n-2}=D^{j,n-2}$ are plotted for $10\leq j\leq n-$ 1, but with the x-axis measured in simulation time units for clarity and illustrative purposes. The y-axis is again on a logarithmic scale. First, it is important to note that for all $j$ and any $n$ (not just the ones shown here), the spectral radii of the update matrices stay below the threshold of $1$, which ensures the convergence of the online smoother as an iterative contraction mapping. Second, the spectral radii, or in this trivial case of one-dimensional latent state space, the absolute value of the update operator $\mathbf{D}^{j,n-2}=D^{j,n-2}$, for every $n$, behaves as an exponentially decreasing function as $j$ decreases down from $n-1$ to 00, which indicates the exponential decrease of the impact region of each new observation by the turbulent dynamics, as discussed in Section 3.2. In contrast, for $n$ values such that $10\leq j\leq n-$ 1 covers an extreme event, the emergence of an instability in the observed signal induces a rapid regime switch and a steep decrease in the spectral radius values, i.e., the influence region of this observation, for when $j$ passes through these intermittent periods, as showed by the steep drops observed at around $t=37.5$ and $t=35$ for each curve covering this period.
@@
Figure 4.2: Panels (a)–(f): Real-time comparison between the filter posterior mean time series (in green) and the one generated from the adaptive-lag online smoother strategy by Theorem 3.1 (in purple). The observed trajectory of $u$ is plotted in panel (a) (in black), while the true trajectory of $v$ is showed in all panels for reference (in blue). Panel (g): Adaptive lag values (measured in time units, i.e., $L_{n}\Delta t$) generated by the algorithm defined in Section 3.4.2, specifically through (3.31)–(3.32). Panel (h): The information gain criterion used to define the adaptive lags through (3.31), for the same time instants as the ones showed in panels (b)–(f), as a function of $j$. These sequences are plotted on a common x-axis for illustration purposes (although their respective $j$’s take different values), for $0\leq j\leq b$, with the x-axis measured in simulation time units. The y-axis is on a logarithmic scale. The threshold value $\delta=10^{-6}$ used in this study is also plotted via a black dashed line for reference. Panel (i): Spectral radii of the update values $D^{j,n-2}$ as a function of $10\leq j\leq n-$ 1 (x-axis is in simulation time units), for the same time instants as the ones displayed in panels (b)–(f). The y-axis is again logarithmically scaled. Panels (h) and (i) share the same legend.

### v2

@@ 4 Applications of the Adaptive-Lag Online Smoother
This section demonstrates the application of the online smoother in addressing three key scientific challenges: (a) state estimation and causality detection, especially in the presence of intermittency and extreme events, (b) evaluating computational and storage efficiency in a high-dimensional Lagrangian data assimilation application, and (c) developing an online parameter estimation algorithm with partial observations.
The simulations for all the numerical experiments in this work were carried out on an AMD RyzenTM 7 5800H mobile CPU of 8 cores, 16 threads, with an average clock speed of 3.2 GHz (turbo boosts up to 4.4 GHz), with a TDP of 45W. In terms of memory, a two-memory-module configuration was used, with 2x16GB DDR4 RAM sticks at 3200MT/s.
@@ 4.1 Detecting Causal Evidence Using the Adaptive-Lag Online Smoother
${\,\rm d}u(t)$ $=\big(-d_{u}u(t)+\gamma u(t)v(t)+F_{u}\big){\,\rm d}t+\sigma_{u}{\,\rm d}W_{u}(t),$ (4.1)
${\,\rm d}v(t)$ $=\big(-d_{v}v(t)-\gamma u(t)^{2}+F_{v}\big){\,\rm d}t+\sigma_{vu}u(t){\,\rm d}W_{u}(t)+\sigma_{v}{\,\rm d}W_{v}(t),$ (4.2)
where all parameters are constants, with $d_{u},d_{v},\sigma_{u},\sigma_{v}>0$, and $\gamma,\sigma_{vu},F_{u},F_{v}\in\mathbb{R}$, while $W_{u}$ and $W_{v}$ are two mutually independent Wiener processes. In (4.1)–(4.2), The variable $u$ represents an observed or resolved mode in the turbulent signal, interacting with the unresolved mode $v$ through quadratic nonlinearities. As $v$ influences the damping of $u$, variations in $v$ cause the time series of $u$ to display intermittency and extreme events. The system (4.1)–(4.2) fits exactly the CGNS framework in (2.1)–(2.2), with $\mathbf{x}=x:=u$ and $\mathbf{y}=y:=v$. The following parameter values are used for this numerical case study:
@@
The parameters are assigned to induce intermittent and extreme events in the observable variable $u$ through the stochastic damping generated by the unobserved variable $v$, with the goal of producing highly non-Gaussian PDFs for both observable and unobserved dynamics. These values are also chosen to address potential issues of observability [12, 90, 91], which is linked to causal detection using the adaptive online smoother. The coupled system in (4.1)–(4.2) loses practical observability when the observed process $u$ provides no information about the unobserved variable $v$. Intuitively, this occurs during phases where $u\approx 0$, rendering $v$ ineffective in contributing to the dynamics of $u$. A numerical integration time step of $\Delta t=0.005$ is used, which also defines the uniform rate of obtaining the observations for the online smoother, with a total simulation time of $T=60$ units.
@@
Figure 4.1 presents various plots comparing the filter and offline smoother posterior estimated states at different levels (this is equivalent to an online smoother which uses $L_{n}=n$ at each $t_{n}=n\Delta t$, where $n$ is the current number of observations; see Section 3.3). Panel (a) displays the trajectory of the observed variable $u$, while Panel (b) displays that for the unobserved variable $v$. For the latter, we also include the posterior means calculated using Theorems 2.2 and 2.3, along with the first two standard deviations away from the mean, represented by accordingly colored shaded regions. The posterior estimated states overall follow the true values with relatively low uncertainty. As expected, the smoother provides a more accurate estimation during the uncertain quiescent periods of the unobserved variable. During the extreme events which are characterized by a high signal-to-noise ratio, both methods lead to smaller uncertainty. Panel (c) plots the information gain of the filter and smoother posterior distributions relative to the prior statistical attractor (or equilibrium distribution) of $v$ on a logarithmically scaled y-axis. Since both the equilibrium (computed numerically using many long simulated trajectories; see Panel (f)), $p_{\text{att}}(v)$, and posterior distributions (at each time instant) are Gaussian, the information gain in Panel (c) is calculated using the signal–dispersion decomposition formula of relative entropy in (3.25). The temporal evolution of relative entropy reveals that the information gain from the smoother posterior Gaussian distribution is almost everywhere larger than that from the filter solution, as is expected. Notably, the information gain corresponding to the onset of the intermittent phases of $u$ shoots up due to the strong signal-to-noise ratios. Note that the information gain using the smoother is more significant than the filter at these phases and the more uncertain quiescent phases due to additional observational information. Panel (d) shows the time-averaged PDF of $u$ generated by its signal over $t\in[20,60]$, alongside the corresponding Gaussian fit, i.e., the normal distribution with the same mean and standard deviation as the truth over the associated period. Due to the intermittent extreme events, we can discern that the induced density is highly non-Gaussian, with an apparent positive skewness and a heavy tail. Panel (e) is the same as (d) but for $v$, which also includes the densities corresponding to the filter and smoother solutions. We can again see that the density corresponding to $v$ displays some non-Gaussian features. Furthermore, as is expected, the PDF of the smoother solution is better equipped to approximate that of the truth compared to the filter one by better resolving the tail behavior and first few ordered moments.
@@
Figure 4.1: Panel (a): Trajectory of the observable variable $u$. Panel (b): True trajectory of the unobserved variable $v$ (in blue), alongside the filter (in green) and offline smoother (in red) posterior mean time series. Their respective first two standard deviations away from the mean state are also plotted through correspondingly colored shaded regions, with the standard deviation in this case being exactly equal to the square root of the respective posterior covariance. The threshold above which $v$ acts as anti-damping to $u$, i.e., the line $y=d_{u}/\gamma=1/6$, is also plotted in a dashed black line. Panel (c): Temporal evolution of the information gain of the filter and smoother posterior Gaussian statistics beyond the statistical attractor of the unobserved variable. This is calculated using the signal–dispersion decomposition of the relative entropy in (3.25). A logarithmic scale is used for the y-axis. Panel (d): Time-averaged PDF of $u$ calculated using the observations over $t\in[20,60]$ (in black), alongside its Gaussian fit density defined by the mean and standard deviation of the signal in the same time period (in dashed magenta). Panel (e): Same as (d) but for the unobserved variable $v$, together with the PDFs corresponding to the filter and smoother posterior mean time series. Panel (f): PDF of the unobserved variable’s prior statistical attractor, $p_{\text{att}}(v)$.
@@
Figure 4.2 focuses on a shorter intermittent period $\big(t\in[32.5,40]\big))$ to assess the performance of the adaptive-lag online smoother using the original sequence of information gains, $\left\{\mathcal{P}\left(p^{j,n}_{\text{updated}},p^{j,n}_{\text{lagged}}\right)\right\}_{R_{n}\leq j\leq n-1}$, with hyperparameters $b=600$ (which corresponds to $b\Delta t=3$ simulation time units) and $\delta=10^{-4}$. For this case study, similar results can be obtained when instead using the LSDev sequence, $\{\sigma^{j,n}\}_{R_{n}\leq j\leq n-1}$, as already discussed in Section 3.4.2, by appropriately adjusting $\delta$ (see also Panel (g) of Figure C.1). Panel (a) shows the signals of $u$ and $v$ over this period, while Panels (b)–(f) illustrate how the online smoother adjusts its state estimation as new observations arrive at five different time instants during this period. The online smoother is compared against the filter solution, which can also produce real-time estimated states with the sequential arrival of observations. We focus on Panels (c) and (d). When using the filter solution, the estimation of $v$ before $u$ reaches the peak of the extreme event contains significant errors. This is because the growth of $v$, once it exceeds the threshold of $d_{u}/\gamma=1/6$ to become anti-damping, triggers the extreme event in $u$ (the cause). However, the extreme event (the effect) does not appear instantly, thus introducing a delay in the onset of these instabilities. This is why at $t=36.50$ (Panel (c)), without seeing the future information from the development of the extreme event, the causal variable $v$ cannot be accurately recovered by the filter. By extension, due to the data from the extreme event in the observed signal simply not being available yet at the current time instant, this, in a sense, contaminates the online smoother state, which faithfully follows the filter one regardless of how long the adaptive lag is. Nevertheless, at $t=37.00$ (Panel (d)), with the observed trajectory currently undergoing the effects of the instability caused by $v$, this allows the adaptive online smoother, which utilizes future information for state estimation unlike the filter one, to correct its previous estimated states based on the value of the calculated adaptive lag, at most for $b\Delta t=3$ time units into the past. This is explicitly seen by the online smoother posterior mean, which now better approximates the true signal compared to its state half a time unit ago, thus demonstrating its property to systematically correct biases when future data become available. This attribute is crucial to help identify the current state, therefore, outperforming the filter solution. This crucial discrepancy between the filter and online smoother estimates reveals that information in the dynamics flows from the current state of the unobserved variable (cause) to the future state of the observed one (effect), highlighting a causal relationship between these states. Panel (g) presents the adaptive lag values generated during this period, from using the sequence of information gains or relative entropies as outlined in (3.32) and (3.31). The lag in the online smoother likewise reflects the aforementioned time delay in the causal relationship, with a peak in adaptive lag values at the onset of the extreme event, i.e., the causal period. These are essential to correct past errors and effectively recover the state of the cause $v$, which triggered the extreme event. In contrast, smaller lag values are needed during the effect period, specifically during the demise of this intermittent event, which is characterized by a high signal-to-noise ratio where both the filter and smoother estimates can effectively capture the relatively deterministic behavior. As such, lower adaptive lag values are required since the “lagged” distribution in (3.28) can yield a quickly decaying information gain function as $j$ decreases from n1n-1 to $R_{n}$ (or to an already small sequence of information gains for these values of $n$, i.e., below the threshold parameter $\delta$ uniformly in $j$), thus leading to small values for $L_{n}$ in (3.32). After all, the lag essentially functions as an interpolation between the past state estimates and a “full” (since $b$ limits how much into the past we can correct) backward-run smoother solution of the online smoother. As such, a close-to-zero lag value signifies that more trust is put in the statistics from the previously calculated online posterior distributions since the observations obtained during this period have tiny impact regions.
Recall that $\left\{\mathcal{P}\left(p^{j,n}_{\text{updated}},p^{j,n}_{\text{lagged}}\right)\right\}_{R_{n}\leq j\leq n-1}$, defined by (3.28)–(3.29), is used to determine the adaptive lags depicted in Panel (g) (via (3.32) and (3.31)). We define the standardization of this sequence in the following manner:
$\frac{\displaystyle\left|\mathcal{P}\left(p^{j,n}_{\text{updated}},p^{j,n}_{\text{lagged}}\right)-\mathcal{P}\left(p^{0,n}_{\text{updated}},p^{0,n}_{\text{lagged}}\right)\right|}{\displaystyle\max_{R_{n}\leq j\leq n-1}\left\{\left|\mathcal{P}\left(p^{j,n}_{\text{updated}},p^{j,n}_{\text{lagged}}\right)-\mathcal{P}\left(p^{0,n}_{\text{updated}},p^{0,n}_{\text{lagged}}\right)\right|\right\}},\quad\forall n\in\mathbb{N},\quad R_{n}\leq j\leq n-1.$ (4.3)
(Other forms of standardization yield similar results, i.e., by dividing by the standard deviation instead of the maximum over $j$.) The standardization of the sequence of its LSDevs, $\{\sigma^{j,n}\}_{R_{n}\leq j\leq n-1}$, is similarly defined. Henceforth, when we refer to the standardized variant of these sequences, we imply the sequence in $j$ and $n$ defined by (4.3).
In Panel (h) of Figure 4.2, the standardized sequence of information gains or relative entropies (as defined in (4.3)) is plotted, on a logarithmic scale, as a function of $n$ (for $n\Delta t\in[32.5,40]$) and of $n\Delta t-j\Delta t$ (for $j\Delta t\in[R_{n}\Delta t,n\Delta t]$). (As a reminder, for the $n$-th observation of $u$, the last index $j\in\{R_{n},\ldots,n-1\}$ for which $\mathcal{P}\left(p^{j,n}_{\text{updated}},p^{j,n}_{\text{lagged}}\right)$ remains below the tolerance value of $\delta=10^{-4}$, is $n-1-L_{n}$.) Of significance here is the observational interval of $n\Delta t\in[35.5,37.5]$, which covers the cause and effect of the extreme event. Up until around $t=37.00$, which is on the cusp of the extreme event’s completion and $u$’s subsequent decline, the cause is fully captured with a persisting standardized information gain throughout the period which generated said extreme event. This explains the increasing and large adaptive lags leading up to the extreme event (i.e., during its generation), depicted in Panel (g). But, as soon as the effect in $u$, generated by $v$, ceases to exist after $t=37.00$, then a causal role reversal is initiated, with $u$ now being the significant factor in the system’s dynamics and in driving $v$’s decrease or damping via the $\gamma u^{2}$ term. As a result, the standardized information gain shows an extremely rapid decay during this regime-switch period, where the high signal-to-noise ratio leads to negligible impact regions for the observations of $u$ on $v$’s online smoother state estimation (into the past, during updates). This has the ulterior outcome of smaller adaptive lag values during this period, as shown in Panel (g), compared to the other periods. Importantly, this behavior persists over the following quiescent period, whenever $n\Delta t-j\Delta t$ bridges over and before $t=37.00$.
@@
Finally, in Panel (i) of Figure 4.2, the spectral radii of the update matrices $\mathbf{D}^{j,n-2}=D^{j,n-2}$ (i.e., their absolute value), are plotted in the same manner as the standardized information gains shown in Panel (h) (though, not standardized). First, it is important to note that for all $j$ and any $n$ the spectral radii of the update matrices stay below the threshold of $1$, which ensures the convergence of the online smoother as an iterative contraction mapping. Second, the spectral radii, or in this trivial case of a one-dimensional latent state space, the absolute value of the update operator $\mathbf{D}^{j,n-2}=D^{j,n-2}$, for every $n$, behaves overall as an exponentially decreasing function as $j$ decreases down from n1n-1 (or as njn-j increases from 0), which indicates the exponential decrease of the impact region of each new observation by the turbulent dynamics, as discussed in Section 3.2.1. As discussed in the prequel and observed in Panel (h), similarly here for $n$ values such that $R_{n}\Delta t\leq j\Delta t\leq(n-1)\Delta t$ covers an extreme event, the emergence of an instability in the observed signal induces a rapid regime switch and an additional steep exponential decrease in the spectral radius values, i.e., the exact influence region of this observation, for when $j$ passes through these intermittent periods. Finally, and most remarkably, is the fact that the information-based criterion we have developed in Section 3.4.2, after standardization, showcases an impressively similar temporal behavior (both in observational (in $n$) and natural (in $j$) time), with slightly differing scales, as the exact metric that defines the observations’ impact or influence region on online smoother state estimation, that being the spectral radii of the update operators (see also Figure 4.5). This is true not just during extreme events, as already discussed, but also during quiescent periods. This provides strong numerical corroboration for our methodology of defining the adaptive lag as an extremely cheap (when compared to the exact method) yet effective approach.
@@
Figure 4.2: Panels (a)–(f): Real-time comparison between the filter posterior mean time series (in green) and the one generated from the adaptive-lag online smoother strategy by Theorem 3.1 (in purple). The observed trajectory of $u$ is plotted in Panel (a) (in black), while the true trajectory of $v$ is showed in all panels for reference (in blue). Panel (g): Adaptive lag values (measured in time units, i.e., $L_{n}\Delta t$) generated by the algorithm defined in Section 3.4.2, specifically through (3.32) and (3.31) (i.e., using the original sequence of relative entropies). Panel (h): Standardized information gain criterion in (4.3) as a function of $n$ (for $n\Delta t\in[32.5,40]$) and of $n\Delta t-j\Delta t$ (for $j\Delta t\in[R_{n}\Delta t,n\Delta t]$). Plotted on a logarithmically scaled colormap. Panel (i): Same as Panel (h) but for the spectral radii of the update values $D^{j,n-2}$, i.e., $|D^{j,n-2}|$. Panels (h) and (i) share the same colorbar, and have their y-axis flipped.
For completeness, since a study into the trade-off between the computational and storage-wise complexity and lower-order pathwise error of the fixed- and adaptive-lag online discrete smoother solutions is provided for the high-dimensional Lagrangian data assimilation problem in Section 4.2 (see Figure 4.4), a similar analysis for this dyad-interaction model is also carried out. It illustrates the computational storage advantages of the adaptive-lag online smoother compared to its fixed-lag variant, for a model defined by intermittent instabilities rather than high dimensionality. See Figure C.1 in Appendix C.